OpenStudy (anonymous):

Help!!!! If the system of equations given below is solved using the substitution method, what is the equation that results from both solving the first equation for x, and substituting into the second equation? x - 7y = 10 2x + 4y = 15 18y = -5 x = 7y + 10 -10y = -5 18x = 145

OpenStudy (anonymous):

does anyone know I am stuck ?

jimthompson5910 (jim_thompson5910):

solve the first equation for x to get x - 7y = 10 x = 10+7y then plug this into the second equation 2x + 4y = 15 2(10+7y) + 4y = 15 20+14y+4y = 15 20+18y = 15 I'll let you finish